A turntable is designed to acquire an angular
velocity of 33.5 rev/s in 0.5 s, starting from
rest.
Find the average angular acceleration of th
turntable during the 0.5 s period.
Answer in units of rad/s
1
Expert's answer
2012-07-19T10:35:01-0400
Let a be the angular acceleration of a turntable. Then

W = aT,

where W is the final angular velocity and T is the acceleration time. So,

a = W/T = (33.5 rev/s) / (0.5 s) = 67 rev/s² = 67·2π rad/s² = 134π rad/s².